OpenStudy (anonymous):

Find the surface area of a sphere if the radius is 2.4 cm

jigglypuff314 (jigglypuff314):

Surface area of a Sphere = 4 (pi) r^2 so given r=2.4 = 4 (pi) (2.4) (2.4) = ...